OpenStudy (anonymous):

What is the length of a cube of volume 0.064 cm3

OpenStudy (anonymous):

lenth=0.64

OpenStudy (anonymous):

seems unlikely

OpenStudy (anonymous):

you need \(\sqrt[3]{.064}\)

OpenStudy (anonymous):

which is a lot more obvious if you write it as \[\sqrt[3]{\frac{64}{1000}}\]

OpenStudy (midhun.madhu1987):